# DexGANGrasp: Dexterous Generative Adversarial Grasping Synthesis for Task-Oriented Manipulation

### v1

@@ Abstract
We introduce DexGanGrasp, a dexterous grasping synthesis method that generates and evaluates grasps with single view in real time. DexGanGrasp comprises a Conditional Generative Adversarial Networks (cGANs)-based DexGenerator to generate dexterous grasps and a discriminator-like DexEvalautor to assess the stability of these grasps. Extensive simulation and real-world expriments showcases the effectiveness of our proposed method, outperforming the baseline FFHNet with an $18.57\%$ higher success rate in real-world evaluation. We further extend DexGanGrasp to DexAfford-Prompt, an open-vocabulary affordance grounding pipeline for dexterous grasping leveraging Multimodal Large Language Models (MLLMs) and Vision Language Models (VLMs), to achieve task-oriented grasping with successful real-world deployments.
@@ I Introduction
Dexterous hands allow for a broader range of solutions to grasp an object. However, they also present the challenge of generating high-quality grasps within a high-dimensional search space. Robotiq-3F hand, the robotic hand used in this work, has in total 18 Degrees of Freedom (DoF) with 6 DoF grasp pose plus 12 DoF joint configurations. To efficiently generate meaningful grasp configurations for dexterous hands, especially from a single view, many approaches adopt learning based methods. These methods may require time-consuming shape completion [7, 8, 9], map visual input to a single unique grasp instead of many [10, 11], predict grasp translation and rotation in a cascaded manner [12], or require additional contact anchor as input [13]. We propose DexGanGrasp , which directly generates high-quality full grasp configurations from only partial point clouds in real-time.
Dexterous hands are designed to mimic human functionality, and humans often grasp objects with a specific purpose in mind, where the method of grasping serves that purpose. For example, a human might grasp a hammer by the handle to use it but grasp the opposite side when handing it over. Different tasks may require completely different ways of grasping the same object. Most existing task-oriented grasping methods are either limited to 2-jaw grippers [14, 15] or rely on human-labeled task grasping datasets [16, 5, 6].
In this work, we propose DexGanGrasp, a method to directly generate diverse dexterous grasps from a single view in real-time. DexGanGrasp consists of a conditional GANs-based DexGenerator which generates various high-quality grasps given random samples from the latent space, and DexEvaluator, which further predicts the grasp stability. Furthermore, we extend DexGanGrasp to task-oriented grasping by introducing DexAfford-Prompt, an open-vocabulary affordance grounding pipeline for dexterous grasping, leveraging MLLMs and VLMs. The user-initiated affordance is fed into a MLLMs together with the detected object name to identify which part of the object supports the task affordance. Subsequently, VLMs segments the part based on the object part name. Finally, grasps generated by DexGanGrasp targeting the segmented part are filtered for execution.
@@
  • We propose a grasp synthesis and evaluation pipeline using conditional GANs for grasping unknown objects with a dexterous hand from a single view in real-time.
@@ II-A Grasp Synthesis
Learning-based grasp synthesis immensely focuses on cVAE [2, 25, 26, 8] and Generative Adversarial Networks (GANs) [7, 27, 28], autoregressive [9] and diffusion models [29].
@@
GANs has been widely studied as a generative model, especially in the field of image synthesis [30, 31, 32], sim2real transfer [33, 34] and so on. Several works adopt GANs for multi-fingered grasp synthesis with RGB-D input data [7, 27]. Because of an integrated shape completion module, their grasping pipeline is slow, taking approximately 8 seconds. In contrast, our model predicts full grasps much faster. Similarly, GANs are employed in [28] to predict a 6D grasp pose along with one of four predefined grasp types from depth images. We apply GANs directly on the point cloud to achieve fast run time while extending it further to task-oriented grasping.
@@ II-B Task-Oriented Grasping
Task-oriented grasping for unknown objects requires robots to grasp object specific parts for specific tasks. Several approaches require a task-labeled grasping dataset [16, 5, 6] or leverage off-the-shelf vision-language models [36, 37].
Graph Convolutional Network (GCN) is utilized in [16] to encode the relationship between objects and tasks. [5] introduces touch code to represent the contacts between the hand and object parts. They further train a model with an attention mechanism to predict the functional grasps. Their model requires manually annotated touch code and full knowledge of the object to predict the grasp. This idea is further extended in [38] for refinement steps to achieve stable and human-like grasps. To increase the grasping dataset quality, [6] introduces a Real-Simulation Annotation System to generate a synthetic dataset with the help of real-world human demonstrations. Our framework, leveraging MLLMs and VLMs, enables task-oriented grasping without needing a dedicated dataset.
[36] takes as input scene point cloud and a task-oriented object part name and outputs a ranked grasps located in this part region. A novel grasp evaluator with Large Language Models (LLMs) input is proposed to achieve zero-shot task-oriented grasping in [37]. Both approaches are applied to a 2-jaw gripper, which limits the manipulability of the following tasks.
@@ III Problem Formulation
This work aims to address the challenge of grasp synthesis for unknown objects with a multi-fingered hand, utilizing only a partial object point cloud $\mathbf{x}\in\mathbb{R}^{N\times 3}$. A grasp $\mathbf{g}\in\mathbb{R}^{d}$ is represented by the 12-DOF hand joint configuration $\boldsymbol{j}\in\mathbb{R}^{15}$ and the 6D palm pose $(\mathbf{R},\mathbf{t})\in SE(3)$. Grasping success for Robotiq-3F hand is defined as its ability to lift an object 20 cm above its resting position without any slippage. Since one object can be grasped in infinite equally successful ways, the underlying relationship between $\mathbf{x}$ and $\mathbf{g}$ is one-to-many instead of one-to-one or many-to-one. Thus, regression or classification-based approaches show limitations in modeling this relationship.
@@ IV Methods
We propose a DexGanGrasp for dexterous grasp synthesis based on conditional GANs, shown in figure 1. The DexGanGrasp consists of a DexGenerator, a DexDiscriminator and a DexEvaluator.
Figure 1: DexGanGrasp consists of a DexGenerator, a DexDiscriminator and a DexEvaluator. DexGenerator takes as input a bps-encoded point cloud as well as a sample $z$ from latent space. It tries to predict various ”fake grasps” $\mathbf{g}=(\mathbf{R},\mathbf{t},\mathbf{j})$, as close as possible to ”real grasps” as ground truth. Meanwhile, DexDiscriminator tries to differentiate between these two types of grasps. During inference, DexEvaluator filters stable grasps out of $\mathbf{g}$ from DexGenerator and the best grasp is chosen for the robot to execute.
@@ IV-A DexGenerator and DexDiscriminator
The DexGenerator takes input of a noise sample from latent space $z\sim p_{z}(z)$ and it generates a grasp distribution $\mathbf{g}\sim p_{g}(z|\mathbf{x})$ conditioned on point cloud $\mathbf{x}$. The DexGenerator is trained to capture the ground truth grasp distribution $p_{r}$ and generate synthetic grasps or ”fake grasps” $\mathbf{g}\sim p_{g}(z|\mathbf{x})$ as realistic as possible. On the other hand, the DexDiscriminator $D$ tries to differentiate between ground truth ”real grasp” $\mathbf{g}\sim p_{r}$ or a ”fake grasp” $\mathbf{g}\sim G(z|\mathbf{x})$ by maximizing $\mathbb{E}_{\mathbf{g}\sim p_{r}(\mathbf{g})}[\log{D(\mathbf{g})}]$ and minimizing $\mathbb{E}_{z\sim p_{z}(z)}[\log{(1-D(G(z|\mathbf{x})))}]$. The training of DexGanGrasp consists of two steps. The first step updates the discriminator weights alone with discriminator loss $L_{D}$ leaving the generator frozen, in equation 1.
@@
The second step will update the generator and discriminator weights together with generator loss $L_{G}$.
For a performant GAN, generating high-fidelity adversarial examples is of high importance in deceiving the discriminator during training. The discriminator loss $L_{D}$ and vanilla generator loss alone are not sufficient to train the DexGanGrasp model, due to the large diversity of the grasp distribution. Therefore, for the generator’s loss $L_{G}$, the adversarial loss $L_{fake\_gen}$ is coupled with a set of distance losses, $L_{dist}$, that guide the model into generating a grasp distribution close to the ground truth.
$L_{dist}=L2_{transl}+L2_{rot}+L2_{joint\_conf}.$ (2)
For the $L_{fake\_gen}$, we label the $\mathbf{g_{gen}}$ generated grasps as real. This is done in order to prevent the DexDiscriminator from easily identifying the generator’s output, since it is already trained with correctly labeled generated grasps during its update stage, defined by $L_{D}$. This, in theory, should create a large gradient when DexDiscriminator is presented with generated grasps, which would backpropagate through the whole network. This optimization stage can also be seen as motivating the generator to synthesize realistic grasps that can’t be easily detected by DexDiscriminator, and also better balancing the learning progress between DexDiscriminator and DexGenerator. This labeling approach is inspired by [39], where the labels are switched stochastically as a method to add noise to the discriminator to balance the adversarial training, as described before. Detailed influence of these tricks are shown in ablation study table IV.
$L_{fake\_gen}=\mathbb{E}_{z\sim p_{z}(z)}[\log{(1-D(G(z|\mathbf{x})))}].$ (3)
So the final loss for the second stage is:
$L_{G}=w_{fake\_gen}L_{fake\_gen}+w_{dist}L_{dist}.$ (4)
@@ IV-B DexEvaluator
$\mathcal{L}_{Eva}=-\left[y\log(f(\mathbf{g}_{pos},\mathbf{x}))+(1-y)\log(f(\mathbf{g}_{neg},\mathbf{x}))\right],$ (5)
@@ IV-C DexGanGrasp at Inference
Afterwards, DexEvaluator predicts scores for all generated grasps $\mathbf{g}$, and the grasp with the best score will be chosen for the robot to execute.
Figure 2: DexAfford-Prompt builds on top of DexGanGrasp as an open-vocabulary affordance grounding pipeline to achieve task-oriented dexterous grasping. Firstly, an RGB image and the question ”What do you see?” are fed to MLLMs to obtain the object name. The detailed prompt is in IV-D. This object name, along with a user-initiated affordance such as ”grab”, is fed into a LLMs to predict the object part name, which affords the task. Further, a VLMs segments the part from image space and projects to 3D space to obtain object part point cloud. A KNN-based filtering function is used to filter out the grasps that do not target the object part. Finally, the grasps are ranked with DexEvaluator for final execution.
@@ IV-D DexAfford-Prompt
We extend DexGanGrasp to a task-oriented grasping pipeline DexAfford-Prompt, described in Algorithm 1. This pipeline is inspired by [14], and we further simplify it by using MLLMs to replace the object detection part consisting of VLMs and a list of object tokens in [14]. Our method takes as input an RGB image and uses MLLMs first to find what is the object on the table. Together with the object name and user-defined affordance, MLLMs predicts the object part name, which corresponds to the affordance described in the following block.
Affordance Prompting with MLLMs Q: Given an image, list [the object name] on the table with simple words such as cup, bottle, book. Please only answer the object’s name. Do not say anything else.” A: [the object name] Q: affordance is: [given by user], please answer [the object part name] from the image based on the given affordance. A: [the object part name]
The predicted object part name is then fed into a VLMs together with an RGB image to obtain the part segmentation. Given part segmentation, which affords the task prompt, we project it to the point cloud to obtain the part point cloud. Afterwards, we apply KNN to filter generated grasp distribution. The filtering cost function is defined based on Euclidean and rotation distance between each grasp and object part point cloud, to ensure a grasp acting on the object part. Finally, this new list of grasps will be ranked with scores produced by DexEvaluator and the robot will execute a top-ranked grasp.
@@
7:  $\mathbf{g_{gen}}\leftarrow\textit{DexGanGrasp}(\mathbf{x})$
@@
9:  $top\_grasp\leftarrow\textit{DexEvaluator}(\mathbf{g_{part}})$
10:  return  $top\_grasp$
@@ IV-E Implementation Details
The code is implemented with PyTorch 2.3.1. All three proposed models are built on top of fully connected residual blocks with skip connections [40]. DexGenerator and DexDiscriminator are both trained with learning rate of $41e-$ 4, batch size of 32 and 60 epochs. DexEvaluator is trained with lr of $41e-$ 4, batch size of 1024 and 30 epochs.
We use ChatGPT 4o [41] as MLLMs for affordance prompting and vlpart [42] as VLMs for part detection. For KNN we set k to 30 grasps.

### v2

@@ Abstract
We introduce DexGanGrasp, a dexterous grasp synthesis method that generates and evaluates grasps with a single view in real-time. DexGanGrasp comprises a Conditional Generative Adversarial Network (cGAN)-based DexGenerator to generate dexterous grasps and a discriminator-like DexEvalautor to assess the stability of these grasps. Extensive simulation and real-world experiments showcase the effectiveness of our proposed method, outperforming the baseline FFHNet with an $18.57\%$ higher success rate in real-world evaluation. To further achieve task-oriented grasping, we extend DexGanGrasp to DexAfford-Prompt, an open-vocabulary affordance grounding pipeline for dexterous grasping leveraging Multimodal Large Language Models (MLLM) and Vision Language Models (VLM) with successful real-world deployments. For the code and data, visit our website.
@@ I Introduction
Dexterous hands allow for a broader range of solutions to grasp an object. However, they also present the challenge of generating high-quality grasps within a high-dimensional search space. Robotiq-3F hand is used in this work. A grasp for Robotiq-3F hand consists of 6 Degrees of Freedom (DoF) grasp pose and 12 DoF joint configurations. To efficiently generate meaningful grasp configurations for dexterous hands, especially from a single view, many approaches adopt learning-based methods. These methods may require time-consuming shape completion [7, 8, 9], map visual input to a single unique grasp instead of many [10, 11], predict grasp translation and rotation in a cascaded manner [12], or require additional contact anchor as input [13].
Dexterous hands are designed to mimic human functionality, and humans often grasp objects with a specific purpose in mind, and the method of grasping serves that purpose. For example, a human might grasp a hammer by the handle to use it but grasp the opposite side when handing it over. Different tasks may require completely different ways of grasping the same object. Most existing task-oriented grasping methods are either limited to 2-jaw grippers [14, 15] or rely on human-labeled task grasping datasets [16, 5, 6].
In this work, we propose DexGanGrasp, a method to directly generate diverse dexterous grasps from a single view in real-time. DexGanGrasp consists of a cGAN-based DexGenerator which generates various high-quality grasps given random samples from Gaussian distributions, and DexEvaluator, which further predicts the grasp stability. Furthermore, we extend DexGanGrasp to task-oriented grasping by introducing DexAfford-Prompt, an open-vocabulary affordance grounding pipeline for dexterous grasping, leveraging MLLM s and VLM s. The user-initiated affordance is fed into a MLLM together with the detected object name to identify which part of the object supports the task affordance. Subsequently, a VLM segments the part based on the object part name. Finally, grasps generated by DexGanGrasp targeting the segmented part are filtered for execution.
@@
  • We propose a grasp synthesis and evaluation pipeline using a cGAN for grasping unknown objects with a dexterous hand from a single view in real-time.
@@ II-A Grasp Synthesis
Learning-based grasp synthesis immensely focuses on cVAE [2, 25, 26, 8] and Generative Adversarial Networks (GAN) [7, 27, 28], autoregressive [9] and diffusion models [29].
@@
GANs have been widely studied as generative models, especially in the field of image synthesis [30, 31, 32], sim2real transfer [33, 34] and so on. Several works adopt GAN s for multi-fingered grasp synthesis with RGB-D input data [7, 27]. Because of an integrated shape completion module, their grasping pipeline is slow, taking approximately 8 seconds. In contrast, our model predicts full grasps much faster. Similarly, GAN s are employed in [28] to predict a 6D grasp pose along with one of four predefined grasp types from depth images. We apply a GAN directly on the point cloud to achieve fast run time while extending it further to task-oriented grasping.
@@ II-B Task-Oriented Grasping
Task-oriented grasping for unknown objects requires robots to grasp object-specific parts for specific tasks [36]. Several approaches require a task-labeled grasping dataset [16, 5, 6] or leverage off-the-shelf vision-language models [37, 38, 39] because of their reasoning capability.
Graph Convolutional Network (GCN) is utilized in [16, 40] to encode the relationship between objects and tasks. [5] introduces touch code to represent the contacts between the hand and object parts. They further train a model with an attention mechanism to predict the functional grasps. Their model requires manually annotated touch code and full knowledge of the object to predict the grasp. This idea is further extended in [41] for refinement steps to achieve stable and human-like grasps. To increase the grasping dataset quality, [6] introduces a Real-Simulation Annotation System to generate a synthetic dataset with the help of real-world human demonstrations. Our framework, inspired by [42], leveraging MLLM s and VLM s, enables task-oriented grasping without needing a dedicated dataset.
[37] takes as input scene point cloud and a task-oriented object part name and outputs a ranked grasps located in this part region. However, they require multiple views to reconstruct the scene, whereas our approach works on a single view. A novel grasp evaluator with Large Language Models (LLM) input is proposed to achieve zero-shot task-oriented grasping in [38]. Both approaches are applied to a 2-jaw gripper, which limits the manipulability of the following tasks. A convex decomposition approach is utilized in [39] to better determine the parts. However, without a dedicated grasp generation module, the grasps mainly end up with top-down grasps, limiting their manipulability afterward. Our proposed DexAfford-Prompt is built on top of a diverse grasp generation module DexGanGrasp for more flexible 6D grasps.
@@ III Problem Formulation
This work aims to address the challenge of grasp synthesis for unknown objects with a multi-fingered hand, utilizing only a partial object point cloud $\mathbf{x}\in\mathbb{R}^{N\times 3}$. A grasp $\mathbf{g}\in\mathbb{R}^{d}$ is represented by the 12-DOF hand joint configuration $\boldsymbol{j}\in\mathbb{R}^{12}$ and the 6D palm pose $(\mathbf{R},\mathbf{t})\in SE(3)$. Grasping success for Robotiq-3F hand is defined as its ability to lift an object 20 cm above its resting position without any slippage. Since one object can be grasped in infinite equally successful ways, the underlying relationship between $\mathbf{x}$ and $\mathbf{g}$ is one-to-many instead of one-to-one or many-to-one. We apply generative models cGAN to tackle this problem.
@@ IV Methods
We propose a DexGanGrasp for dexterous grasp synthesis based on cGAN, shown in figure 1. The DexGanGrasp consists of a DexGenerator, a DexDiscriminator and a DexEvaluator.
Figure 1: DexGanGrasp consists of a DexGenerator, a DexDiscriminator and a DexEvaluator. DexGenerator takes as input a bps-encoded point cloud as well as a sample $z$ from latent space. It tries to predict various ”fake” grasps $\mathbf{g_{gen}}=(\mathbf{R},\mathbf{t},\mathbf{j})$, as close as possible to the positive ground truth ”real” grasps $\mathbf{g}_{pos}$. Meanwhile, DexDiscriminator tries to differentiate between these two types of grasps. During inference, DexEvaluator filters stable grasps out of $\mathbf{g}$ from DexGenerator and the best grasp is chosen for the robot to execute.
@@ IV-A DexGenerator and DexDiscriminator
The DexGenerator $G$ takes the input of a noise sample from latent space $z\sim p_{z}(z)$ defined as a Standard Normal Gaussian $N(0,1)$, and generates a synthetic grasp distribution $\mathbf{g}\sim p_{g}(z|\mathbf{x})$ conditioned on point cloud $\mathbf{x}$. The DexGenerator is trained to capture the ground truth grasp distribution $p_{r}$ and generate synthetic or ”fake” generated grasps $\mathbf{g_{gen}}\sim G(z|\mathbf{x})$ as realistically as possible. On the other hand, the DexDiscriminator $D$ tries to differentiate between the positive ground truth ”real” grasp $\mathbf{g}_{pos}\sim p_{r}$ or a ”fake” grasp $\mathbf{g_{gen}}$ by maximizing $\mathbb{E}_{\mathbf{g}_{pos}\sim p_{r}(\mathbf{g}_{pos})}[\log{D(\mathbf{g}_{pos})}]$ and minimizing $\mathbb{E}_{z\sim p_{z}(z)}[\log{(1-D(G(z|\mathbf{x})))}]$.
The training of DexGanGrasp consists of two steps. The first step updates the discriminator weights alone with discriminator loss $\mathcal{L}_{D}$ leaving the generator frozen, in eq. 1.
@@
The second step will update the generator and discriminator weights together with generator loss $\mathcal{L}_{G}$.
The discriminator loss $\mathcal{L}_{D}$ and vanilla generator loss alone are not sufficient to train the DexGanGrasp model, due to the large diversity of the grasp distribution. Therefore, we define the generator’s loss $\mathcal{L}_{G}$ as a combination of the adversarial loss $\mathcal{L}_{fake\_gen}$ and $\mathcal{L}_{dist}$, a weighted sum of $L2$ distance losses from translation, rotation, and joint configuration. $\mathcal{L}_{dist}$ guides the model into generating a grasp distribution close to the ground truth. The weights are determined manually to put all the losses on the same scale.
$\mathcal{L}_{G}$ $=L_{fake\_gen}+L_{dist}$ (2)
$\mathcal{L}_{fake\_gen}$ $=\mathbb{E}_{z\sim p_{z}(z)}[\log{(1-D(\mathbf{g_{gen}}))}]$ (3)
$\mathcal{L}_{dist}$ $=w_{t}L_{transl}+w_{r}L_{rot}+w_{j}L_{joint}$ (4)
@@ IV-B DexEvaluator
$\mathcal{L}_{Eva}=-\left[y\log(f(\mathbf{g}_{pos},\mathbf{x}))+(1-y)\log(f(\mathbf{g}_{neg},\mathbf{x}))\right]$ (5)
@@ IV-C DexGanGrasp at Inference
Afterward, DexEvaluator predicts scores for all the generated grasps $\mathbf{g_{gen}}$, and the grasp with the best score will be chosen for the robot to execute.
Figure 2: DexAfford-Prompt builds on top of DexGanGrasp as an open-vocabulary affordance grounding pipeline to achieve task-oriented dexterous grasping. Firstly, an RGB image and the question ”What do you see?” are fed to ChatGPT 4o (MLLM[43] to obtain the object name. The detailed prompt is in IV-D. This object name, along with a user-initiated affordance such as ”grab”, is fed into ChatGPT 4 (LLM[43] to predict the object part name, which affords the task. Further,VLPart (VLM[44] segments the part from image space and projects to 3D space to obtain object part point cloud. A KNN-based filtering function is used to filter out the grasps that do not target the object part. Finally, the grasps are ranked with DexEvaluator for final execution.
@@ IV-D DexAfford-Prompt
We extend DexGanGrasp to a task-oriented grasping pipeline DexAfford-Prompt, described in Algorithm 1. This pipeline is inspired by [14], and we further simplify it by using MLLM s to replace the object detection part consisting of a VLM and a list of object tokens in [14]. Our method takes as input an RGB image and uses the MLLM first to find what is the object on the table. Together with the object name and user-defined affordance, the LLM predicts the object part name, which corresponds to the affordance described in the following block.
Affordance Prompting with MLLM/LLM Q: Given an image, list the object name on the table with simple words such as cup, bottle, and book. Only answer the object’s name. Don’t say anything else.” A: [the object name] Q: affordance is: [given by user], please answer [the object part name] from the image based on the given affordance. A: [the object part name]
The predicted object part name is then fed into the VLM together with an RGB image to obtain the part segmentation. Given part segmentation, which affords the task prompt, we project it to the point cloud to obtain the part point cloud. Afterwards, we apply KNN to filter generated grasp distribution. The filtering cost function is defined based on Euclidean and rotation distance between each grasp and object part point cloud, to ensure a grasp acting on the object part. Finally, this new list of grasps will be ranked with scores produced by DexEvaluator and the robot will execute a top-ranked grasp.
@@
7:  $\mathbf{g_{gen}}\leftarrow\textit{DexGenerator}(\mathbf{z},\mathbf{x})$
@@
9:  $filtered\_grasps\leftarrow\textit{DexEvaluator}(\mathbf{g_{part}})$
10:  $top\_grasp\leftarrow select\_best(filtered\_grasps)$
11:  return  $top\_grasp$
@@ IV-E Implementation Details
The code is implemented with PyTorch 2.3.1. All three proposed models are built on top of fully connected residual blocks with skip connections [45]. In $\mathcal{L}_{fake\_gen}$ we label all the $\mathbf{g_{gen}}$ generated grasps as real. This is done in order to prevent the DexDiscriminator from easily identifying the generator’s output since it is already trained with correctly labeled generated grasps during its update stage, defined by $\mathcal{L}_{D}$. This, in theory, should create a large gradient when DexDiscriminator is presented with generated grasps, which would backpropagate through the whole network. This labeling approach is inspired by [46], where the labels are switched stochastically as a method to add noise to the discriminator to balance the adversarial training. DexGenerator and DexDiscriminator are both trained with learning rate of $41e-$ 4, batch size of 32 and 60 epochs. DexEvaluator is trained with lr of $41e-$ 4, batch size of 1024 and 30 epochs.
We use ChatGPT 4o [43] as MLLM for object detection, ChatGPT 4 [43] as LLM for affordance prompting, and VLPart [44] as VLM for part detection. For KNN, we set k to 30 grasps.